OpenStudy (anonymous):

How would I write y+2=7/8(x-3) in standard form?

OpenStudy (anonymous):

y+2=7/8(x-3) y+2=0.875(x-3) y+2=0.875x-2.625 -2 -2 y=0.875x-4.625

OpenStudy (anonymous):

y=0.875x-4.625 +0.875x +0.875x y+0.875x=-4.625